OpenStudy (anonymous):

Identify the 16th term of a geometric sequence where a1 = 4 and a8 = -8,748

hartnn (hartnn):

do you know the formula for n'th term of geometric sequence ?

OpenStudy (anonymous):

an= a1 x r ^n-1

hartnn (hartnn):

correct! now put n= 8 there.... a8 is -8748 n=8, a1 =4 only 'r' is unknown, can you find it ?

OpenStudy (anonymous):

how do you find the r ?

OpenStudy (anonymous):

can you please help me find the 16th term though

hartnn (hartnn):

-8748 = 4 8 r^7 once you get the value of common ratio 'r' finding the 16th term is just one step! :)

hartnn (hartnn):

*** -8748 = 4 *r^7 can you find r from here ?

OpenStudy (anonymous):

no can you help me please

hartnn (hartnn):

-8748 = 4 *r^7 divide both sides by 4 what u get ?

OpenStudy (anonymous):

is it -2187 = xx ^n-7

hartnn (hartnn):

-2187 = r^7 now use the calculator to get 7th root of 2187

OpenStudy (anonymous):

is ir 46.7653718

hartnn (hartnn):

no, its simply -3 because -3^7 = -2187

hartnn (hartnn):

now we have r=-3 and we need 16th term, so we put n=16 and plug these in the n'th term equation, an= a1 x r ^n-1

hartnn (hartnn):

a16 = 4 (-3)^15 = ... ?

OpenStudy (anonymous):

is it -8748

hartnn (hartnn):

no, just use a calculator to calculate 4*(-3)^15

OpenStudy (anonymous):

yea thank you i got it